# The Physics of Local Optimization in Complex Disordered Systems

### v2

@@ Abstract
In summary, using exact GS algorithms, we study the effectiveness of local subsystem solvers and relate it to critical threshold ZED physics [18] of system-spanning avalanche-like vanishing-energy excitations. In short-range spin-glasses, away from these special critical thresholds, the response of the system to local perturbations to the GS problem decays rapidly. The difference between local minimization and its full global counterpart becomes significant primarily near the critical coupling threshold values. Away from critical threshold values, local GS solvers achieve increased accuracy. Importantly, the associated error rate of the local solvers becomes independent of the system size [see Fig. 2 (a)] for large systems and decreases rapidly with the subsystem size. Local solvers can thus become increasingly effective for large $L$ where a local solution of a subsystem that is a very small fraction $(L_{\rm sub}/L)^{d}$ of the large system may correctly predict local properties of the large system with negligible error. Simple functions link the average error rate of local solvers to their (subsystem) size as well as to deviations of the coupling constants from their critical threshold value, and changes of the critical coupling of one bond as a result of tuning another distant bond to its critical threshold value. Our analysis reveals that various length-dependent quantities appear to display algebraic scaling just like the scale-free distributions of avalanche size [18], while for ferromagnetic phases we observe an exponential decay of errors with subsystem size (see also End Matter and [28], Sec. LABEL:sec:shifted_gaussian). These behaviors may rationalize the success of more sophisticated local solvers in spin-glasses [24, 25, 26] and other systems [1, 2], and might be broadly applicable in systems not constrained by the “Overlap Gap Property” [42]. For the all-to-all SK model, on the other hand, the error rates always depend on the size of the full system. These differences lead us to conjecture a classification of possible levels of local hardness (Ref. [28], Sec. LABEL:sec:level_hardness). Local solvers are found to be efficient for the considered lattices irrespective of whether the global problems are of P (2D) or NP (3D) complexity. The occurrence of rare events in the form of bonds attaining their critical couplings explains why local solvers cannot in general guarantee to find exact GSs for the full systems.
We briefly speculate on broad implications and extensions of our results regarding the effectiveness of local optimization. Efficient and often highly accurate local minimization is of immense utility in replacing more taxing global (“batch”) gradient descent methods by inherently fast sequential local stochastic gradient descent-type approaches across numerous problems including high-dimensional/large-data machine learning type tasks in many-body physics, materials science, community detection, and other domains [43, 44, 45, 46, 47, 48]. Local plastic variants of the backpropagation algorithm used in machine learning might emulate efficient biological neural computing [49, 50]. Other notable aspects of local minimization and optimized structure in artificial and real neural networks have been further investigated [51, 52]. Locality is of paramount importance in numerous problems in many-body physics [3, 4, 5, 6, 7] where the degrees of freedom may also be continuous. Replacing Eq. (1) by $H=-\sum_{ij}J_{ij}s_{i}s_{j}+u\sum_{i}(s_{i}^{2}-1)^{2}$ with $u\gg 1$ (Eq. (1) corresponds to $u\to\infty$) for continuous $s_{i}\in\mathbb{R}$ suggests that the phenomena that we found might extend to continuous spin systems of sufficiently large quartic interaction strength $u$. In a specific continuous spin variant, that of the spherical model [53, 54], the locality that we described here is altered (Ref. [28], Sec. LABEL:sec:solvable). In future work, we will also further explore viable relations to previously unreported criticality (Ref. [28], Sec. LABEL:sec:sg_criticality). Ultimately, the physics of local optimization may underlie Nature’s enigmatic efficiency, shaping computation across diverse realms.
@@
Acknowledgments. We thank Mark Goh for a discussion in which he alerted us to [42]. Z.N. and M.W. are grateful for support from TU Chemnitz through the Visiting Scholar Program. Z.N. is further grateful for support through the Washington University Seeding Projects for Enabling Excellence & Distinction (SPEED) and the Leverhulme-Peierls Senior Researcher Professorship at Oxford via a Leverhulme Trust International Professorship Grant No. LIP2020-014 (S. L. Sondhi). G.O. gratefully acknowledges support from the Institute for Advanced Study.

### v3

@@ Abstract
In summary, using exact GS algorithms, we study the effectiveness of divide-and-conquer type local subsystem solvers and relate it to critical threshold ZED physics [shen2023universal] of system-spanning avalanche-like vanishing-energy excitations. In short-range spin-glasses, away from critical thresholds, the response of the system GS to local perturbations decays rapidly. Generally, the difference between local and global minimization only becomes significant near critical coupling threshold values. Away from these critical values, local GS solvers achieve increased accuracy. Importantly, the associated error rate of the local solvers becomes independent of the system size [see Fig. 2 (a)] for large systems and decreases rapidly with the subsystem size. Local solvers may thus become increasingly effective for large $L$. Our analysis reveals that various length-dependent quantities display algebraic scaling just like the scale-free distributions of avalanche size [shen2023universal], while for ferromagnetic phases we observe an exponential decay of errors with subsystem size (see also End Matter and [SM], Sec. LABEL:sec:shifted_gaussian). These behaviors may rationalize the success of more sophisticated solvers in spin-glasses [wang_patch-planting_2017, RL_local, RL_local1] and other systems [Imagenet, Alphafold], and might be broadly applicable in systems not constrained by the “Overlap Gap Property” [OGP]. For the all-to-all SK model, on the other hand, the error rates always depend on the size of the full system. These differences lead us to conjecture a classification of possible levels of local hardness (Ref. [SM], Sec. LABEL:sec:level_hardness). Local solvers are found to be efficient for the studied lattices irrespective of whether the global problems are of P (2D) or NP (3D) complexity. The occurrence of rare events of bonds being close to their critical couplings explains why local solvers cannot, in general, guarantee to find exact GSs of the full systems. This mechanism which is reminiscent of chaos in spin glasses [fernandez2013temperature, baity2021temperature, katzgraber_temperature_2007] is generic for the classes of systems we study. It is instructive to briefly relate the local-hardness perspective to the metastate and chaotic-size-dependence framework [Newmanbook]: there, one asks whether local observables have an effectively “concentrated” or more “dispersed” distribution over thermodynamic states. In our setting, large stability $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ corresponds to a computable proxy for a robust metastate-like concentration of local observables of low LSBS error while small stability reflects competition between nearly degenerate states.
While these considerations were focused on the hyper-local task of predicting the spins on both ends of a single bond, it is possible to stitch together the results of such calculations to candidate ground states of the full system. In the End Matter and Ref. [SM] (Sec. Contraction solver based on LSBS), we outline an LSBS-based solver that uses the subsystem critical thresholds $J_{\mathrm{c},ij}^{\rm sub}$ to identify “well-determined” bonds, i.e., bonds for which $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ is large and the LSBS error rate is exponentially small, cf. Eq. (3). This identification enables the contraction of the system into a smaller number of cohesive large-scale elements (or “communities” [Girvan_Newman2002, Fortunato2010, Comm_detection_local, Comm_detection_local']) before solving for the GS. This extension is here presented as proof-of-concept demonstration of how local hardness can guide algorithm design, and further detailed study is required to gauge its full potential. Generalized to settings that include external fields and multi-body interactions (Ref. [SM], Sec. LABEL:sec:multibody), our framework becomes applicable to other important NP-hard optimization problems, such as the Minimum Vertex Cover problem (see End Matter and Ref. [SM], Sec. LABEL:sec:mvc_ising). Locality further underlies numerous many-body physics problems [Kohn-near, Cluster1, Cluster2, Cluster3, Cluster4] where degrees of freedom may also be continuous; replacing Eq. (1) by $H=-\sum_{ij}J_{ij}s_{i}s_{j}+u\sum_{i}(s_{i}^{2}-1)^{2}$ with $u\gg 1$ (Eq. (1) corresponds to $u\to\infty$) for continuous $s_{i}\in\mathbb{R}$ suggests that our findings extend to continuous systems of sufficiently large quartic interaction strength $u$. In a continuous spin variant defined by a global constraint– the spherical model [Berlin+Kac, kosterlitz_spherical_1977]– the locality that we described here is altered (Ref. [SM], Sec. LABEL:sec:solvable). Another interesting application area for the ideas outlined here lies in machine learning, with the potential of replacing slow global gradient methods by fast sequential local stochastic gradient descent-type approaches [Harju1997, Comm_detection_local, Comm_detection_local', Ruder2017, review_ML_SGD, MatDis]. Local plastic variants of the backpropagation algorithm used in machine learning might emulate efficient biological neural computing [Song2020, Goodfellow]. Other notable aspects of local minimization and optimized structure appear in biological and artificial neural networks [NN_modular, modular_brain]. We view our LSBS-based contraction solver as a starting point for future investigations of local-hardness–guided heuristics. Dividing large scale problems into smaller easier-to-solve fragments may also be beneficial for quantum computing schemes for which smaller scale systems might suffer less from decoherence. Extensions of the GS problem analyzed here, including products between distant spins (see Ref. [SM], Sec. LABEL:sec:distant_spins) to finite temperature correlation functions may provide illuminating connections between our results and finite size scaling [FS1, FS2, FS3, FS4]. In future work, we will further detail relations to previously unreported criticality (Ref. [SM], Sec. LABEL:sec:sg_criticality). The physics of local optimization may underlie Nature’s efficiency, shaping computation across diverse realms.
@@
Acknowledgments. We thank Mark Goh for a discussion in which he alerted us to [OGP]. Z.N. and M.W. are grateful for support from TU Chemnitz through the Visiting Scholar Program. Z.N. is further grateful for support through the Washington University Seeding Projects for Enabling Excellence & Distinction (SPEED) and the Leverhulme-Peierls Senior Researcher Professorship at Oxford via a Leverhulme Trust International Professorship Grant No. LIP2020-014 (S. L. Sondhi). G.O. gratefully acknowledges support from the Institute for Advanced Study.
End Matter
Here, we provide numerical evidence for Eqs. (3) and (4), the existence of a transition in the LSBS accuracy as the distribution of couplings is shifted (so as make the system more ferromagnetic and easier to solve), and we discuss beyond nearest-neighbor correlations found by the LSBS.
Error rates for deviations from the critical threshold
Illustrating the applicability of the analysis in the main text, Fig. 4 demonstrates that the LSBS error indeed decays with increasing $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ in a manner that adheres to the stretched exponential drop of Eq. (3).
Figure 4: Error rate $\mathcal{E}_{ij}$ as a function of $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ for (a) a 2D (square lattice), $L=1024$ system with $L_{\rm sub}=8$, $16$, $32$, $64$ and (b) a 3D (cubic) system of $L=12$ with $L_{\rm sub}=4$, $10$. Point percentages denote cumulative probabilities (i.e., fraction of instances) for values of $|J_{\mathrm{c},ij}^{\rm sub}-J_{ij}|$ lower than their abscissa.
Figure 5: Average change of critical threshold values ($|\Delta{J_{\mathrm{c},ij}}|]$) of bonds a distance of $r$ along a Cartesian direction from -$\infty$ to $\infty$ in (a) 2D ($L=128$) and (b) 3D ($L=12$) systems. Dashed: Eq. (4) with parameters in Table 1. As is more evident in 2D, when $r={\cal{O}}(L)$, the average $|\Delta J_{\mathrm{c},ij}|]$ is bounded from above by these ultra-local forms with the fitted $\ell_{c}$ being of the order of the lattice constant.
.1 Distance dependence of critical threshold values
As discussed in the main text, the critical threshold of a given bond is largely determined only by the couplings of other bonds that are very close to it. When we tuned the central bonds of $L=128$ square and $L=12$ cubic lattices to their critical values, the respective critical threshold changes ($\Delta J_{\mathrm{c},ij}$) of all other bonds at a distance $r$ away conformed to Eq. (4) with $\ell_{c}$ of the order of the lattice constant, see Table 1 (with the fit of Eq. (4) becoming an upper bound for the largest distances in the 2D system). This is illustrated in Fig. 5. In Ref. [SM], Sec. LABEL:sec:correl_critical_couplings (see Fig. LABEL:fig:visualization_Jc_change therein), we further provide a calculated real-space visualization of such a typical rapid drop of the change in the critical coupling away from the origin. We further observe a smaller effect on distant bonds for smaller ZEDs.
Figure 6: For an asymmetric bond distribution, as $\overline{J}$ increases the system transitions from a spin-glass phase to a disordered ferromagnetic phase [melchert_scaling_2009]. We observe how the fitting coefficient $\kappa$ changes as we adjust $\overline{J}$. A larger value of $\kappa$ implies that the error rate of the subsystem solver decays faster, with a crossover to an exponential decay beyond the threshold $1/\overline{J}=0.96$, where the systems enters the ferromagnetic phase [melchert_scaling_2009].
[…]
Figure 7: Schematic of the contraction solver. (a) Well-determined bonds ($|J_{ij}-J_{\mathrm{c},ij}|\geq\vartheta$, where $\vartheta$ is a manually chosen threshold) form connected components that are contracted to effective sites (shown as larger nodes), inheriting all interactions between the component and the rest of the system. (b) The relative spin orientations (indicated by ++ and - signs) of sites within each connected component are locked in by the well-determined bonds. The GS of the full system can then be reconstructed from the GS of the contracted system by propagating the component orientations. Different colors represent distinct connected components.
Contraction solver based on LSBS
The LSBS provides local predictions for individual bond spin products. Here we outline how these local predictions can be combined into a contraction solver that approximates the full GS spin configuration. For each bond, we compute its subsystem critical threshold $J_{\mathrm{c},ij}$ and identify well-determined bonds satisfying $|J_{ij}-J_{\mathrm{c},ij}|\geq\vartheta$, where $\vartheta$ is a chosen threshold. These bonds, being far from their critical thresholds, have high local prediction accuracy. Well-determined bonds form connected components (or “communities” [Girvan_Newman2002, Fortunato2010, Comm_detection_local, Comm_detection_local']) within which relative spin orientations are reliably fixed. Each such component is contracted to a single effective site, inheriting all interactions between the component and the rest of the system. The contracted system is solved exactly, and the original spin configuration is reconstructed by propagating the component orientations. This contraction procedure is illustrated schematically in Fig. 7. For appropriate choices of $\vartheta$, this approach yields configurations very close to the exact GS while substantially reducing the effective problem size. Detailed schematics, algorithmic steps, and performance analysis are provided in the Supplemental Material (Ref. [SM], Sec. Contraction solver based on LSBS).
Minimum Vertex Cover as an Ising problem
As a representative NP-hard optimization problem beyond spin glasses, we also examined the Minimum Vertex Cover (MVC) on random planar graphs generated by Delaunay triangulation. Mapping a vertex $i$ being in the cover to $\sigma_{i}=+1$ and otherwise $\sigma_{i}=-1$, the cost function can be written as
$H_{\text{MVC}}=A\sum_{i\in V}\frac{1+\sigma_{i}}{2}+B\sum_{\langle ij\rangle\in E}\frac{1-\sigma_{i}}{2}\frac{1-\sigma_{j}}{2},$ (5)
with couplings $J_{ij}=-B/4$ on edges $\langle ij\rangle\in E$ and fields $h_{i}=-A/2+(B/4)\deg(i)$ (for $A=1$, $B=2$). Using LSBS on subsystems of topological radius $r$ around a central site, we predict its relative spin orientation in the MVC GS and compare to the exact solution, finding a power-law-like decay of the LSBS error rate with subsystem size (Fig. 8) with exponent $\kappa\approx 0.8$, close to the spin-glass case. This suggests that our locality-based hardness picture applies more broadly to combinatorial optimization problems.
Figure 8: Minimum Vertex Cover (MVC) as an Ising problem on random planar graphs. LSBS error rate for the central spin in the MVC mapping, of the random planar graph shown in the inset, as a function of subsystem size, showing a power-law-like decay with exponent $\kappa\approx 0.8$. (Lower left inset) Random planar graph generated from $n=10^{4}$ points by Delaunay triangulation, with sites colored by topological distance $r$ from the geometric center.